how do you express 1698 as a roman numeral?

Found 2 solutions by benni1013, CoolRekt:
Answer by benni1013(206)
Answer by CoolRekt(53)

MDCXCVIII = 1698
M = 1000
D = 500
C = 100
XC = 90
V = 5
III = 3
M + D + C + XC + V + III = 1000 + 500 + 100 + 90 + 5 + 3 = MDCXCVIII = 1698
CoolRekt